A gaseous mixture contains 406.0 Torr of H2(g), 347.9 Torr of N2 g), and 83.1 Torrof Ar(g). Calculate the mole fraction, X, of each of these gases. Number Number Number
Answer #1

find mol frac of each gas

total P = 406+347.9+83.1 = 837 torr

now

x-H2 = P-H2/Total P = 406/837 = 0.48506 torr

x-N2 = P-N2/Total P = 347.9/837 = 0.41565 torr

x-Ar = P-Ar/Total P = 83.1/837 = 0.09928 torr